Any help with this system of linear equation would be appreciated.
In a triangle, the sum of the interior angles is 180 degrees. If angle A is three times as large as angle B, and angles B and C sum to 90 degrees, what is the measure of each angle?

Answer by Earlsdon(6294)

From the problem description, you can write:
A = 3B "...angle A is three times as large as angle B,..."
B+C = 90 "...angles B and C sum to 90 degrees,..."
A+B+C = 180 Substitute B+C = 90
A+90 = 180 Subtract 90 from both sides.
A = 90 degrees. Substitute A = 3B.
3B = 90 Divide both sides by 3.
B = 30 degrees.
B+C = 90 Substitute B = 30.
30+C = 90 Subtract 30 from both sides.
C = 60 degrees.
-----------------
Angle A = 90 degrees.
Angle B = 30 degrees.
Angle C = 60 degrees.